Looks like some sort of DoS
MAJOR Closed Broadband and Ethernet
STATUS
Closed
AFFECTED
Broadband and Ethernet
STARTED
Mar 29, 07:00 PM (8½ months ago)
CLOSED
Apr 04, 06:00 PM (8¼ months ago)
REFERENCE
2512 / AA2512
INFORMATION
  • INITIAL
    8½ months ago by Adrian

    Over the last few days we have seen several short attacks on parts of our network. Staff are monitoring and taking action. We don't usually put a lot of detail regarding DoS attacks. In some cases services may be briefly unavailable and routing may change. Sorry for any inconvenience.

  • UPDATE
    8½ months ago by Adrian

    Overall traffic for broadband customers does not seem to be that impacted, but clearly something not quite right, and we are still looking in to it.

  • UPDATE
    8½ months ago by Adrian

    Things are not quite perfect but mostly looking back to normal now.

  • UPDATE
    8½ months ago by Adrian

    It has died down a lot, with some action from the ops team, but not entirely sorted you. Bear with us...

  • UPDATE
    8½ months ago by Adrian

    Sorry for delay posting, we are seeing the problem again this afternoon.

  • UPDATE
    8½ months ago by Adrian

    Mostly broadband traffic not impacted - but services in Maidenhead are being hit, including VoIP and web servers.

  • UPDATE
    8½ months ago by Adrian

    Again some issues, and again a slight delay posting a status as effort on ensuring service working first, sorry about that. Some really odd things this last day or so...

  • UPDATE
    8½ months ago by Adrian

    Once again, it is worth explaining that the majority of traffic levels are sensible but this is causing some brief annoyances - it almost seems to be randomly targeting customers. We are keeping a close eye on the issues.

  • UPDATE
    8½ months ago by Adrian

    Obviously we are a bit limited with what we can say regarding attacks. At present the web servers are having issues, sorry for any inconvenience.

  • UPDATE
    8½ months ago by Adrian

    OK, something more severe right now, looking in to it

  • UPDATE
    8½ months ago by Adrian

    This seems to be impacting all broadband.

  • UPDATE
    8½ months ago by Adrian

    That was weird, something very odd there and may have been one of our processes that caused an issue this evening, we are reviewing now.

  • UPDATE
    8½ months ago by Andrew

    Problems are happening again this afternoon - staff have been on call all weekend working behind the scenes on this.

  • UPDATE
    8½ months ago by Adrian

    It took slightly longer chasing the attack this evening. Some transit disruption.

  • UPDATE
    8¼ months ago by Adrian

    Sorry, reporting every blip, even if only a few seconds or even if once in a day, is probably not helping and may feed the attacker. I can say it is being works on at many levels and seems to be diminishing. Sorry for inconvenience.

  • UPDATE
    8¼ months ago by Adrian

    We are seeing routing issues to Maidenhead at present (web sites, voip). This seems to be a result of work we are doing now to address some of the issues of the last few days.

  • UPDATE
    8¼ months ago by Adrian

    This looks stable now.

  • UPDATE
    8¼ months ago by Adrian

    Still ongoing this evening, and we are improving the responses and have even added some pro-active measures. We are working on it, and having to take much more severe measures than ever before, but we are not really able to go in to a huge amount of detail, sorry. Please bear with us.

  • UPDATE
    8¼ months ago by Adrian

    No issues late last night - we're continuing to work on this, obviously.

  • UPDATE
    8¼ months ago by Adrian

    We're keeping this ticket open a little longer - nothing for nearly 48 hours, but staff are still monitoring closely.

  • RESOLUTION
    8¼ months ago by Adrian

    Ongoing work is underway following our experience from these events.

  • Closed